Calculate the market equilibrium if the equation for a market demand curve is Qd = 100 - 0,5P and the equation for market supply curve is Qs = -20 + P.
1
Expert's answer
2020-03-03T09:36:51-0500

In equilibrium, Qd = Qs

or, 100 - 0,5P = -20 + P

120=1,5P

or, P = 80

So, Equilibrium quantity , Q = Qd = Qs = 100 - 0,5х80 = 60 [аs, P = 80 unit.]

Answer: Equilibrium price is 80 unit and quantity is 60 unit.